# Phase transitions in debt recycling

### v1

@@ 1 Introduction
The plan of the paper is as follows. In Section 2, we briefly review the existing literature on mortgages and risks of default, and strategies for equity release. In Section 3, we set up the model and discuss the relevant variables and dynamical equations. In Section 4 we briefly summarize our results for the phase diagram and the main findings from analytical considerations and numerical simulations. In Section 5 we describe the mathematical methods we employ to lead to the analytical solution of the model. In Section 6 we report an extended discussion of results, including parameter dependence and different scenarios. Finally, in Section 7 we offer some concluding remarks and outlook for future research. The Appendix is devoted to technical details.
@@ 2 Related Works
In this section we analyse the main literature related to (i) modelling and data analysis of mortgage defaults and (ii) assessment of debt recyclying and other equity extraction strategies.
@@ 2.1 Models and Analyses of Mortgage Default
The life-cycle model illustrated in [7] shows that households face income and house-price risk when buying homes with mortgages. Combining recourse mortgages and loan-to-valie (LTV) limits effectively reduces default rates, and increases housing demand. The effect of monthly payment size on mortgage default is studied by Fuster et al. in [8]. They find that payment reductions substantially lower mortgage defaults, even for borrowers that are deeply underwater: for instance, cutting the required payment in half reduces the delinquency hazard by about $55\%$.
@@ 3 Model Setup
In this section, we introduce a discrete-time model for the evolution of mortgage and equity values via a debt-recyclying strategy incorporating the following update mechanisms: (i) monthly loan repayments, accounting for the principal and interest components of each monthly mortgage payment arranged with the lender, (ii) fluctuations of the house market that affect the property value, (iii) dynamics of equity investment, as a fraction of their equity is invested via different routes.
@@
We assume that the usable equity at time $t$ – the amount that that can be used to back an investment – is a fraction of the equity owned
@@
The question is therefore how to characterize $t^{\star}$, the first hitting time of the stochastic process onto the absorbing boundaries $E=0$ or $M=0$, and consequently which boundary is reached first.
In the following section, we are going to define the probability distributions and parameter assumptions of our model. Later, we are going to summarize the results for the hitting time of the average processes $\langle E_{t}\rangle$ and $\langle M_{t}\rangle$, where the averages are taken over the realizations of the random variables involved.
@@ 3.1 Assumptions
    The scheduled repayment to the bank $\pi_{t}$, introduced in Eq. (4), is assumed to be non-negative ($\pi_{t}>0$) at every time step $t$, indicating consistent payments to the bank. However, our model allows for the rare event of a few installments of the repayment being skipped (with probability $q$). Otherwise, with probability $1-q$ each installments has value $\pi^{\star}$.
@@
    $s\in[-4\%,+4\%]$: average of the fluctuations in the house market value.
@@
    $\pi^{*}$: value of the monthly installment;
@@ 4 Summary of Results
In Table 1, we summarize the various outcomes of the strategy in different regions of the phase diagram, and the typical process dynamics in each case.
In Fig. 3 we present the phase diagram of the debt recycling strategy with the following parameters: average initial equity $\langle E_{0}\rangle=\$30,000$, i.e., the amount that the homeowner paid at the beginning of the mortgage agreement, or equivalently the value of the portion of the house that the borrower owns totally from the beginning, free from any liabilities or mortgages; average initial mortgage $\langle M_{0}\rangle=\$300,000$, i.e., the part of the property’s purchase price that is financed through the loan; value of the installment of the repayment to the bank $\pi^{*}=\$3,000$; $q=1\%$ chance of skipping an installment; LTV ratio $\ell=0.5$; risk factor of the investment $\mu=0.5$.
@@
Figure 4: Another representation of the phase diagram in Fig. 3, with focus on the strong (dark orange) and weak (light orange) success of the strategy. In cyan, the contour line of $t^{*}=t_{\text{ no recycling}}=\langle M_{0}\rangle/\pi^{*}=25$ years for our choice of parameters $\langle M_{0}\rangle=\$300,000$ and $\pi^{*}=\$3,000$.
@@ 6.1 Simulations of trajectories for mortgage balance and equity
(a) Equity over time. The average process, Eq. (30), is illustrated by the solid green line; some simulations of the process, Eq. (10), are shown in black; the standard deviation band, calculated using $N=10^{5}$ simulations, is displayed in gray.
(b) Mortgage over time. The average process, Eq. (31), is illustrated by the solid red line; some simulations of the process, Eq. (10), are shown in black; the standard deviation band, calculated using $N=10^{5}$ simulations, is represented in gray. The vertical line at $t^{\star}=11.4$ signals the moment the average mortgage balance hits the absorbing wall at zero.
@@
(a) Equity over time. The average process, Eq. (30), is illustrated by the solid green line; some simulations of the process, Eq. (10), are shown in black; the standard deviation band, calculated using $N=10^{5}$ simulations, is displayed in gray. The vertical line at $t^{*}=3.2$ signals the moment the average equity hits the barrier at zero.
(b) Mortgage over time. The average process, Eq. (31), is illustrated by the solid red line; some simulations of the process, Eq. (10), are shown in black; the standard deviation band, calculated using $N=10^{5}$ simulations, is displayed in gray.
@@ 6.2 Phase Diagrams and dependence on parameters
In Sec. 4 we presented a phase diagram with the following parameters. Average initial equity $\langle E_{0}\rangle=\$30,000$; average initial mortgage $\langle M_{0}\rangle=\$300,000$; value of the installment of the repayment to the bank $\pi^{*}=\$3,000$; $q=1\%$ chance of skipping an installment; $\ell=0.5$; $\mu=0.5$. Indeed, under unfavorable investment parameters ($p\sim 0$ and $s\sim-4\%$), the house is entirely mortgaged, and the strategy defaults; with favorable parameters ($p\sim 1$ and $s\sim+4\%$), the strategy is strongly successful. We now analyze and present in Fig. 7 how the phase diagram changes for a wider range of initial conditions (average equity $\langle E_{0}\rangle$, average mortgage $\langle M_{0}\rangle$, average installment of the repayment to the bank $\langle\pi\rangle$), but keeping the same value of the product of the LTV ratio and risk factor $\ell\mu$ [^3] The official data available on the Australian Bureau of Statistics (ABS), updated to December 2023, states that the mean price of residential dwellings is \$$\$933,800$. We chose to reference Australian data because this strategy is already in use there.:
@@
The comparison between Case $1a$ and $1b$ shows the dependence on the initial conditions $\langle E_{0}\rangle$ and $\langle M_{0}\rangle$ separately, with fixed ratio $c_{1}=1$; the comparison between Case $2a$ and $2b$ shows the dependence on the initial conditions $\langle E_{0}\rangle$ and $\langle M_{0}\rangle$ separately, with fixed ratio $c_{1}=10$; the comparison between Case $2a$ and $3a$ shows the dependence on the value of repayment installments $\langle\pi\rangle$.
@@
Comparing Case $1a$ with Case $1b$, and Case $2a$ with Case $2b$, where the difference is in the initial conditions of equity and mortgage, but with the mortgage-to-equity ratio $c_{1}$ kept constant, we find that the areas in the phase diagram for the three different possible outcomes of the strategy remain largely the same, with the purple area slightly bigger when the overall house value is higher. What changes is the speed at which the processes come to a definite outcome. In the orange area, where the strategy is successful, an increase in the total value of the house $H$ ($H=E+M$) corresponds to a slightly longer time required to fully repay it, converting its entire value into equity. Meanwhile, again in the case of a higher house value, the time to reach failure in the area where the strategy fails (in purple) is shorter. Where the strategy is successful, a higher house value means there is more capital to be managed and repaid. This results in a longer time period needed to transform all mortgage into equity, even though the strategy ultimately succeeds. In contrast, in scenarios where the strategy fails, the financial strain of maintaining large mortgage payments against a failing investment strategy becomes unsustainable more quickly, leading to a faster default. These observations imply that while higher property values can enhance the potential gains from successful debt recycling, they also increase the risks and consequences of failure, requiring a more extended commitment period.
@@
Comparing Case $1a$ and $1b$ with Case $2a$ and $2b$, which differ in the initial mortgage-to-equity ratio $c_{1}$, we notice that a higher $c_{1}$ ratio, indicating a higher percentage of the mortgage yet to be repaid relative to the already owned equity, results in an expanded default outcome space, extending well into the IV quadrant. This expansion suggests that even with a good chance of investment gains ($p>0.5$), if the average market value trend is negative ($s<0$), the strategy on average still defaults.
Comparing Case $2a$ with Case $3a$, where the difference lies in the scheduled monthly repayment to the bank, we see that the phase diagram shifts, enlarging the success zone as the value of the scheduled repayment increases.
@@
Therefore, the most influential parameters are the mortgage-to-equity initial ratio, rather than the individual values of mortgage and equity themselves (meaning the strategy’s outcome does not depend strongly on whether we are dealing with a low or high-value house) and the monthly repayments.
@@
Now, we analyze the phase diagrams for different values of $\ell$ and $\mu$ (see Fig.8) . As showed in Sec. 5, the product $\ell\mu\in[0,1]$ is the critical factor influencing the amplitude of the fluctuations of $E_{t}$ and $M_{t}$.
@@
Comparing Case $4a$ with Case $4b$, we observe that increasing the value of the product $\ell\mu$ yields two effects: it reduces the default zone in the IV quadrant and accelerates the strategy outcome. With higher $\ell\mu$, the system becomes more sensitive to changes in parameters $p$ and $s$. Small variations in these parameters can lead to significant changes in the outcome, as indicated by the steeper gradients in Case $4b$. While higher $\ell\mu$ values reduce the risk of default and accelerate success, they also imply that rapid changes of the outcome of the strategy are more likely. Additionally, one may note that only low values of $\ell\mu$ may lead to a weakly successful strategy in the II quadrant, as outlined in Tab. 1.
The straightforward comparison between Case $5a$ and Case $5b$, where only the risk factor $\mu$ increases, leads to qualitatively similar conclusions. We can either increase the LTV ratio parameter, the risk parameter, or both simultaneously for an amplified effect: the result is a faster process, and in the IV quadrant, default is shifted back in favor of success. This effect is amplified as the total value of the house decreases.
@@ 6.2.1 Velocity of the outcome
Taking Fig. 3 as a reference, we travel along the vertical line at $p=0.4$ from bottom to top, increasing the value of $s$. At first, we are deeply into the default (purple) region, but increasing $s$ the time it takes for the average equity to hit zero increases. This is in clear agreement with the housing market becoming progressively more favorable, as also evidenced by the first hitting root of the average equity curves in Fig. 2. At $s=-1.28\%$, we cross the purple-orange boundary in the phase diagram in Fig. 3, meaning that the strategy turns into a (weakly) successful one (light orange). The time required to gain ownership of the house indeed turns out to be longer than without any strategy – just paying back the lender on a monthly basis (cyan line). The time required to repay the debt without applying the strategy is computed assuming a constant repayment amount $\pi^{*}$ to the bank at each time step. With an initial mortgage $M_{0}$ of \$300,000 and $\pi^{*}$ set at \$3,000, the required repayment period is $100$ time steps (or $25$ years). For visual clarity, in the phase diagram we have limited the maximum color intensity to $t=400$ quarters; instances where debt recycling is effective beyond $100$ years are here represented at $t=400$, as well. At $s=0\%$ we further cross the orange-gray boundary and enter the phase of permanent re-mortgaging.

### v2

@@ Abstract
[…]
[4]organization=London Mathematical Laboratory, addressline=8 Margravine Gardens, city=London, postcode=WC 8RH, country=UK
@@ 1 Introduction
The plan of the paper is as follows. In Section 2, we briefly review the existing literature on mortgages and risks of default, and strategies for equity release. In Section 3, we set up the model and discuss the relevant variables and dynamical equations. In Section 4 we briefly summarize our results for the phase diagram and the main findings from analytical considerations and numerical simulations. In Section 5 we describe the mathematical methods we employ to lead to the analytical solution of the model. In Section 6 we report an extended discussion of results, including parameter dependence and different scenarios. In Section 7, we outline the policy implications derived from our model. Finally, in Section 8 we offer some concluding remarks and outlook for future research. The Appendices are devoted to technical derivations, including the exact calculation of process fluctuations, and a discussion of the robustness of conclusions drawn from the outcome of the average process.
@@ 2 Related Works
In this section we analyze the main literature related to (i) modelling and data analysis of mortgage defaults and (ii) assessment of debt recycling and other equity extraction strategies.
@@ 2.1 Models and Analyses of Mortgage Default
The life-cycle model illustrated in [7] shows that households face income and house-price risk when buying homes with mortgages. Combining recourse mortgages and loan-to-value (LTV) limits effectively reduces default rates, and increases housing demand. The effect of monthly payment size on mortgage default is studied by Fuster et al. in [8]. They find that payment reductions substantially lower mortgage defaults, even for borrowers that are deeply underwater: for instance, cutting the required payment in half reduces the delinquency hazard by about $55\%$.
@@ 3 Model Setup
In this section, we introduce a discrete-time model for the evolution of mortgage and equity values via a debt-recyclying strategy incorporating the following update mechanisms: (i) quarterly loan repayments, accounting for the principal and interest components of each monthly mortgage payment arranged with the lender, (ii) fluctuations of the house market that affect the property value, (iii) dynamics of equity investment, as a fraction of their equity is invested via different routes.
@@
We assume that the usable equity at time $t$ – the amount that can be used to back an investment – is a fraction of the equity owned
@@
The question is therefore how to characterize $t^{\star}$, the first hitting time of the stochastic process onto the absorbing boundaries $E=0$ or $M=0$, and consequently which boundary is reached first. First-passage problems for coupled stochastic processes are notoriously hard to tackle exactly [23]. To make the problem more tractable, we are going to focus on the average processes $\langle E_{t}\rangle$ and $\langle M_{t}\rangle$, where the averages are taken over the realizations of the random variables involved. We will also study analytically the fluctuations around the average processes and discuss the robustness of the outcome with respect to such fluctuations.
In the following section, we are going to define the probability distributions and parameter assumptions of our model. Later, we are going to summarize the results for the hitting time of the average processes $\langle E_{t}\rangle$ and $\langle M_{t}\rangle$.
@@ 3.1 Assumptions
    The scheduled repayment to the bank $\pi_{t}$, introduced in Eq. (4), is assumed to be non-negative ($\pi_{t}$geq 00) at every time step $t$, indicating consistent payments to the bank. However, our model allows for the rare event of a few installments of the repayment being skipped (with probability $q$). Otherwise, with probability $1-q$ each installments has value $\pi^{\star}$.
@@
    $s\in[-4\%,+4\%]$: average of the fluctuations in the house market value;
@@
    $\pi^{\star}$: value of the monthly installment;
@@ 4 Summary of Results
In the following Table, we summarize the various outcomes of the strategy in different regions of the phase diagram, and the typical process dynamics in each case.
@@
In Fig. 3 we present the phase diagram of the debt recycling strategy with the following parameters: average initial equity $\langle E_{0}\rangle=\$30,000$, i.e., the amount that the homeowner paid at the beginning of the mortgage agreement, or equivalently the value of the portion of the house that the borrower owns totally from the beginning, free from any liabilities or mortgages; average initial mortgage $\langle M_{0}\rangle=\$300,000$, i.e., the part of the property’s purchase price that is financed through the loan; value of the installment of the repayment to the bank $\pi^{\star}=\$3,000$; $q=1\%$ chance of skipping an installment; LTV ratio $\ell=0.5$; risk factor of the investment $\mu=0.5$.
@@
Figure 4: Another representation of the phase diagram in Fig. 3, with focus on the strong (dark orange) and weak (light orange) success of the strategy. In cyan, the contour line of $t^{*}=t_{\text{ no recycling}}=\langle M_{0}\rangle/\pi^{\star}=25$ years for our choice of parameters $\langle M_{0}\rangle=\$300,000$ and $\pi^{\star}=\$3,000$.
@@ 5 Methods
We refer to B for the calculation of the fluctuations (variances) of the processes $E_{t}$ and $M_{t}$ at all times. The fluctuations will be shown to be fundamentally incapable of changing the conclusions drawn on the basis of the average processes alone, and summarized in the table and phase diagrams.
@@ 6.1 Simulations of trajectories for mortgage balance and equity
(a) Equity over time. The average process, Eq. (30), is illustrated by the solid green line; some simulations of the process, Eq. (10) with a standard deviation $\phi=1\%$ of the fluctuations in the house market value, are shown in black; the standard deviation band, calculated using $N=10^{5}$ simulations, is displayed in gray.
(b) Mortgage over time. The average process, Eq. (31), is illustrated by the solid red line; some simulations of the process, Eq. (10) with a standard deviation $\phi=1\%$ of the fluctuations in the house market value, are shown in black; the standard deviation band, calculated using $N=10^{5}$ simulations, is represented in gray. The vertical line at $t^{\star}=11.4$ signals the moment the average mortgage balance hits the absorbing wall at zero.
@@
(a) Equity over time. The average process, Eq. (30), is illustrated by the solid green line; some simulations of the process, Eq. (10) with a standard deviation $\phi=1\%$ of the fluctuations in the house market value, are shown in black; the standard deviation band, calculated using $N=10^{5}$ simulations, is displayed in gray. The vertical line at $t^{*}=3.2$ signals the moment the average equity hits the barrier at zero.
(b) Mortgage over time. The average process, Eq. (31), is illustrated by the solid red line; some simulations of the process, Eq. (10) with a standard deviation $\phi=1\%$ of the fluctuations in the house market value, are shown in black; the standard deviation band, calculated using $N=10^{5}$ simulations, is displayed in gray.
@@ 6.2 Phase Diagrams and dependence on parameters
In Sec. 4 we presented a phase diagram with the following parameters. Average initial equity $\langle E_{0}\rangle=\$30,000$; average initial mortgage $\langle M_{0}\rangle=\$300,000$; value of the installment of the repayment to the bank $\pi^{\star}=\$3,000$; $q=1\%$ chance of skipping an installment; $\ell=0.5$; $\mu=0.5$. Of course, under unfavorable investment parameters ($p\sim 0$ and $s\sim-4\%$), the house is entirely mortgaged, and the strategy defaults; with favorable parameters ($p\sim 1$ and $s\sim+4\%$), the strategy is strongly successful. We now analyze and present in Fig. 7 how the phase diagram changes for a wider range of initial conditions (average equity $\langle E_{0}\rangle$, average mortgage $\langle M_{0}\rangle$, average installment of the repayment to the bank $\langle\pi\rangle$), but keeping the same value of the product of the LTV ratio and risk factor $\ell\mu$ [^3] The official data available on the Australian Bureau of Statistics (ABS), updated to December 2023, states that the mean price of residential dwellings is \$$\$933,800$. We chose to reference Australian data because this strategy is already in use there.:
@@
6.2.1 Impact of property value with initial mortgage and equity ratio
We first analyze the impact of the property value on the strategy outcomes. We provide two examples (Case $1a$-$1b$ and Case $2a$-$2b$) where we compare - ceteris paribus - debt recycling strategy outcomes for properties with lower ($a$ panels) or higher ($b$ panels).
@@ 6.2.1 Impact of property value with initial mortgage and equity ratio
The comparison between Case $1a$ and $1b$ shows the dependence on the initial conditions $\langle E_{0}\rangle$ and $\langle M_{0}\rangle$ separately, with fixed ratio $c_{1}=1$; the comparison between Case $2a$ and $2b$ shows the dependence on the initial conditions $\langle E_{0}\rangle$ and $\langle M_{0}\rangle$ separately, with fixed ratio $c_{1}=10$.
By comparing these cases, we see the effect of different property values (low vs. high) with a fixed mortgage-to-equity ratio ($c_{1}$). Although the overall structure of success and failure regions remains similar, there are subtle but important differences:
  • 1.
@@
    Time to outcome: For higher property values (e.g., Case $1b$ and Case $2b$ compared with Case $1a$ and Case $2a$ respectively), the strategy requires more time to reach either success or failure. In the orange area, where the strategy is successful, an increase in the total value of the house $H$ ($H=E+M$) corresponds to a slightly longer time required to fully repay it, converting its entire value into equity. Meanwhile, again in the case of a higher house value, the time to reach failure in the area where the strategy fails (in purple) is shorter. Where the strategy is successful, a higher house value means there is more capital to be managed and repaid. This results in a longer time period needed to transform all mortgage into equity, even though the strategy ultimately succeeds. In contrast, in scenarios where the strategy fails, the financial strain of maintaining large mortgage payments against a failing investment strategy becomes unsustainable more quickly, leading to a faster default.
  • 2.
    Expansion of failure region: The purple zone, where debt recycling fails, slightly expands with increased property values. This suggests that while debt recycling has potential rewards for high-value properties, it also entails heightened risks, as these properties amplify both gains and losses. In scenarios with unfavorable market conditions, this expanded failure zone underscores the potential for accelerated default due to increased financial strain.
These observations imply that while higher property values can enhance the potential gains from successful debt recycling, they also increase the risks and consequences of failure, requiring a more extended commitment period.
6.2.2 Impact of Mortgage-to-Equity Ratio
The comparison between cases $1a-2a$ and $1b-2b$ highlights the effect of changing the mortgage-to-equity ratio $c_{1}$ for low and high value properties, respectively.
  • 1.
@@ 6.2.2 Impact of Mortgage-to-Equity Ratio
    Risk of Default in High $c_{1}$ Scenarios: A higher mortgage-to-equity ratio ($c_{1}$), as seen in Cases $2a$ and $2b$ (independently of the initial house value, high vs. low respectively), results in an expanded default outcome space, extending well into the IV quadrant. This suggests that even with a good chance of investment gains ($p>0.5$), if the average market value trend is negative ($s<0$), the strategy on average still defaults when the initial debt burden is significantly higher relative to the property’s equity.
  • 2.
    Implications for debt recycling feasibility: As we will discuss in Sec. 7, this sensitivity to the $c_{1}$ ratio implies that debt recycling strategies are more likely to be successful when borrowers have a lower mortgage-to-equity ratio. Borrowers starting with a high debt relative to equity should be more cautious, as this increases the likelihood of default, even under relatively favorable investment probabilities.
6.2.3 Effect of Increased Repayment Size
The comparison between Case $2a$ and $3a$ shows, ceteris paribus, the dependence on the value of repayment installments $\langle\pi\rangle$, increased in these scenarios from \$3,000 to \$10,000.
  • 1.
    Enlarged success region: Increasing the repayment significantly widens the orange (success) zone even in the III quadrant where $s<0$ and $p<0.5$, indicating that higher repayments improve the likelihood of successful debt recycling. This is because larger repayments reduce the outstanding mortgage more quickly, thereby increasing equity and reducing debt exposure over time.
  • 2.
    Reduced default risk: The purple (failure) zone shrinks when the repayment is increased, as the faster reduction of debt mitigates the risk of default under a range of market conditions. Financially, this suggests that borrowers with the capacity for higher repayments are more likely to sustain a debt recycling strategy successfully, as they can more readily manage and reduce debt exposure.
@@ 6.2.3 Effect of Increased Repayment Size
To summarize, the most influential parameters are the mortgage-to-equity initial ratio, rather than the individual values of mortgage and equity themselves – meaning the strategy’s outcome does not depend strongly on whether we are dealing with a low or high-value house – and the size of the monthly repayments. The implications for policymakers and consideration on feasibility for borrowers will be discussed in Sec. 7.
@@
6.2.4 Dependence on risk factor and LTV ratio
Now, we analyze the phase diagrams for different values of $\ell$ and $\mu$. As showed in Sec. 5, the product $\ell\mu\in[0,1]$ is the critical factor influencing the amplitude of the fluctuations of $E_{t}$ and $M_{t}$.
@@ 6.2.4 Dependence on risk factor and LTV ratio
Comparing Case $4a$ with Case $4b$, we observe that increasing the value of the product $\ell\mu$ yields two effects: it reduces the default zone in the IV quadrant and accelerates the strategy outcome. With higher $\ell\mu$, the system becomes more sensitive to changes in parameters $p$ and $s$. Small variations in these parameters can lead to significant changes in the outcome, as indicated by the steeper gradients in Case $4b$. While higher $\ell\mu$ values reduce the risk of default and accelerate success, they also imply that rapid changes of the outcome of the strategy are more likely.
Additionally, one may note that only low values of $\ell\mu$ may lead to a weakly successful strategy in the II quadrant, as outlined in Tab. 1.
The straightforward comparison between Case $5a$ and Case $5b$, where only the risk factor $\mu$ increases, leads to qualitatively similar conclusions.
We can either increase the LTV ratio parameter, the risk parameter, or both simultaneously for an amplified effect: the result is a faster process, and in the IV quadrant, default is shifted back in favor of success. This effect is amplified as the total value of the house decreases.
@@ 6.2.5 Velocity of the outcome
Taking Fig. 3 as a reference, we travel along the vertical line at $p=0.4$ from bottom to top, increasing the value of $s$. At first, we are deeply into the default (purple) region, but increasing $s$ the time it takes for the average equity to hit zero increases. This is in clear agreement with the housing market becoming progressively more favorable, as also evidenced by the first hitting root of the average equity curves in Fig. 2. At $s=-1.28\%$, we cross the purple-orange boundary in the phase diagram in Fig. 3, meaning that the strategy turns into a (weakly) successful one (light orange). The time required to gain ownership of the house indeed turns out to be longer than without any strategy – just paying back the lender on a monthly basis (cyan line). The time required to repay the debt without applying the strategy is computed assuming a constant repayment amount $\pi^{\star}$ to the bank at each time step. With an initial mortgage $M_{0}$ of \$300,000 and $\pi^{\star}$ set at \$3,000, the required repayment period is $100$ time steps (or $25$ years). For visual clarity, in the phase diagram we have limited the maximum color intensity to $t=400$ quarters; instances where debt recycling is effective beyond $100$ years are here represented at $t=400$, as well. At $s=0\%$ we further cross the orange-gray boundary and enter the phase of permanent re-mortgaging.